In a recent​ year, person A and person B had two concert​ tours, and together they generated $192.9 million in ticket sales. If person B took in ​$22.3 million less than person​ A, how many millions of dollars did each tour​ generate?
Answer by addingup(3677)

A+A-22.3 = 192.9 Add A's on left and add 22.3 on both sides
2A = 215.2 Divide both sides by 2
A = 107.6 and
B = 107.6-22.3 = 85.3
J